OpenStudy (melissa_something):

Log (x^2-7x+12): Find the domain?

OpenStudy (melissa_something):

I know you gotta set it to 0 :(

OpenStudy (anonymous):

I can give you an example. But not the answer. That would be cheating.

OpenStudy (melissa_something):

Lol, the thing is, I already have the answer I'm just not sure how :(

OpenStudy (melissa_something):

I got up to (x-3)(x-4) but that's it.

OpenStudy (anonymous):

See if this helps you! Given the rational function: f(x)=7x / x^2-16 a. The domain When checking domain look for terms with sqrt and zeros in the denominator which would make function undefined. Since there are no sqrts in given equation, check for zeros in the denominator. x^2-16=0 (x+4)(x-4)=0 x=-4 and x=4 domain:{x|x≠-4, x≠4}, or (-∞,-4) U (-4, 4) U (4, ∞)

OpenStudy (anonymous):

Does it help you?

OpenStudy (melissa_something):

I'm looking at it but you gotta set it to >0 in my example. It's the same format but I dont even know how to set it in interval form

OpenStudy (anonymous):

@mathsciencehistory

OpenStudy (anonymous):

wwhy?

OpenStudy (anonymous):

Could you help with this?

OpenStudy (anonymous):

i could try but why you pick me...just random ?

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

@texaschic101

OpenStudy (melissa_something):

how odd. the log doesnt matter btw

OpenStudy (anonymous):

OH OKAY I REMEMBER WHAT TO DO? f(x)= log(x^2-7x+6) x^2-7x<=6 This is not allowed. x^2-7x-6<=0 Does this help?

OpenStudy (anonymous):

So..... f(x)= log(x^2-7x+12) x^2-7x<=12 This is not allowed. x^2-7x-12<=0

OpenStudy (anonymous):

In your case^

OpenStudy (melissa_something):

Lemme see!!!

OpenStudy (anonymous):

Okay take your time.

OpenStudy (melissa_something):

Got it!

OpenStudy (anonymous):

Cool! If you're done close the question, and if you need more help close the question make a new one and tag me. Have a great day ma'am!!!

OpenStudy (anonymous):

Oh BTW you get a medal for effort!!!